A data processing unit has four rows of terminals with six terminals in each row. A second unit has nine rows of terminals with five terminals in each row. The number of additional terminals in the second unit is
Answer by JBarnum(2146)

first has 4 rows with 6 terminals/row thats 6*4=24 terminals
second has 9 rows with 5 terminals/row thats 5*9=45 terminals
45-24=21
there are 21 more terminals in the second unit